OpenStudy (anonymous):

x_x I need help with factoring. [MEDAL :3 thanks in advance]

OpenStudy (anonymous):

\[100x ^{2}-180xy+81y\]

OpenStudy (anonymous):

D; I can't find the GCF</3

hartnn (hartnn):

hint : 100 *81 =10*10*9*9 = (10*9) * (10*9)

hartnn (hartnn):

now, can you factor ?

OpenStudy (anonymous):

D; Um...

OpenStudy (anonymous):

I think so, let me try [DRAGONBALL Z YAY! cx]

hartnn (hartnn):

sure, take your time :) i'll be here if you need further help . [DBZ YAY High %!]

hartnn (hartnn):

*[High 5!]

OpenStudy (anonymous):

OHKAY. I think I have it? ;/ (10x-9y^2)?

hartnn (hartnn):

firstly, i believe the Q is \(100x ^{2}-180xy+81y^2\) secondly, i believe that you meant \((10x-9y)^2\) which is indeed correct! :)

OpenStudy (anonymous):

Yes that. xD Hehe! HURRAH! Thankyou very much :D

hartnn (hartnn):

welcome very much :D